OpenStudy (nottim):

Let f(x,y) = 8e^{2xy}. Evaluate the following partial derivatives at (1,-1). As an option, answers may be written using e and an exponent instead of entering large or small numerical values. Find fxx, fxy and fyy

OpenStudy (anonymous):

Okay, first we need to find the first "general" partial with respect to x, as well as respect to y. What would that be?

OpenStudy (nottim):

Is it fx=2y(8e^(2xy))?

OpenStudy (anonymous):

Yes, that's correct for the first partial with respect to x. What is first partial with respect to y? Then you can take the second derivatives you need to and evaluate.

OpenStudy (nottim):

alright, so fy=2x(8e^2y)?

OpenStudy (anonymous):

Right.. now if you take the partial with respect to x and derive it again, you find: \[f_{xx}(x,y)\] As well as deriving the partial with respect to why according to y again: \[f_{yy}(x,y)\]

OpenStudy (nottim):

Would it be fx=fxx?

OpenStudy (nottim):

And fy=fyy?

OpenStudy (anonymous):

No, \[f_x(x,y) = \frac{\delta}{\delta x}\] \[f_{xx}(x,y) = \frac{\delta}{\delta x} \frac{\delta f}{\delta x} = \frac{\delta^2f}{\delta x^2} = \frac{\delta}{\delta x} \left( f_x(x,y)\right)\]

OpenStudy (anonymous):

In other words, fxx(x,y) is analogous to your f''(x), except it has two variables instead of one(and one of them is "constant" when taking the derivative)

OpenStudy (nottim):

I only know of your final statement there

OpenStudy (nottim):

thanks to khan's academy

OpenStudy (anonymous):

A first-order partial of a two variable function, f(x,y), is the derivative of one of those variables(say, x), assuming that the other variable remains constant(y). In comparison, the second-order partial(fxx(x,y) is the second derivative of one variable(x) assuming the other variable is constant(y).

OpenStudy (nottim):

uh

OpenStudy (nottim):

Wait wait You lost me now

OpenStudy (anonymous):

I'm basically putting into my own words the definition of partial derivatives, both first and second derivatives. If you are confused, how would you describe it so I can see where you might be getting confused.

OpenStudy (nottim):

Alright, sos heres what i did so far

OpenStudy (nottim):

wrong one

OpenStudy (nottim):

this one

OpenStudy (anonymous):

Those are all correct answers as far as I can tell. All that's left is the mixed partial: \[f_{xy}(x,y)\] Now, if everything is nice, this also happens to equal: \[f_{yx}(x,y)\] But we won't worry about it. Instead we can just take the partial we already have with respect to x: \[f_x(x,y)\] And take the partial of *that* with respect to y. As some extra practice, you can take the partial with respect to y and take the derivative of that with respect to x and see if you come up with the same answer both ways.

OpenStudy (nottim):

Ok. I understand. Thank you very much, and also to those who have taken a look at this question.

OpenStudy (nottim):

oops. what did i do wrong?

OpenStudy (anonymous):

Your first step should look like: \[ f_{xy}=2(8e^{2xy}) + 2y(8e^{2xy})(2x) \]

OpenStudy (nottim):

why is the end 2x?

OpenStudy (anonymous):

\[f_x(x,y) = 2y8e^{2xy}, u = 2y, v = 8e^{2xy}, u_y = 2, v_y = (2x)8e^{2xy}\]\[ f_{xy}(x,y) = uv_y + vu_y = 2(8e^{2xy}) + 2y(8e^{2xy})(2x)\]

OpenStudy (nottim):

i opposite it was opposite variables? that what im interpreting ti as from my notes

OpenStudy (nottim):

What about fyy?

OpenStudy (nottim):

@heril ?

OpenStudy (nottim):

Oh

OpenStudy (nottim):

An arthimetic error. Sorry to disturb you...

OpenStudy (nottim):

I figured it out